OpenStudy (secret-ninja):

Not really a question, but just wondering if there are any other JW's on OS. If you are a JW then you'll know what I mean.

OpenStudy (compassionate):

Jehovah's Witnesses? Justice Warriors? SOCIAL Justice Warriors?

TheSmartOne (thesmartone):

\(\begin{array}{lr} \,\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \overset{\bcancel{\;\,}}\rhd \overset{\cancel{\;\,}}\lhd\\ \underset{\widehat\qquad}\blacktriangle\\ \Large \color{#CD5555}{\overline{\quad}} \end{matrix} \right)}}\normalsize \\ \begin{array}{lcr} /&\color{blue}\bowtie&\backslash\\ |&\scriptsize^\circ&|\\ |&\scriptsize^\circ&|\\ \end{array} & \begin{array}{l} \large \text{You have posted in the}\\ \large\text{wong area. Please read }\\ \large\text{the Code of Conduct.} \end{array} \end{array}\)

TheSmartOne (thesmartone):

Got it from: @e.mccormick http://openstudy.com/users/e.mccormick#/updates/5237f559e4b0b9d6b2d6dad6

OpenStudy (confluxepic):

Cool.

OpenStudy (confluxepic):

I am going to try and make something like that.

OpenStudy (e.mccormick):

Hehe, the displeased man has resurfaced.

TheSmartOne (thesmartone):

ikr :)

TheSmartOne (thesmartone):

I am going to use it now :)